# MGDA Converges under Generalized Smoothness, Provably

### v4

@@ 1 Introduction
There have been a variety of emerging applications of multi-objective optimization (MOO), such as online advertising [28], autonomous driving [18], and reinforcement learning [34]. Mathematically, the MOO problem takes the following formulation.
@@
where $K$ is the total number of objectives and $f_{k}(x)$ is the $k$-objective function given model parameters $x$. Under the stochastic setting, $f_{k}(x)=\mathbb{E}_{s}[f_{k}(x;s)]$, where $s$ denotes data samples. This problem is challenging due to the gradient conflict that some objectives with larger gradients dominate the update direction at the sacrifice of significant performance degeneration on the less-fortune objectives with smaller gradients. A variety of MOO-based methods have been proposed to mitigate this conflict and find a more balanced solution among all objectives. In particular, the multiple gradient descent algorithm (MGDA) [12] aims to find a conflict-avoidant (CA) update direction that maximizes the minimal improvement among all objectives and converges to a Pareto stationary point at which there is no common descent direction for all objective functions. This idea then inspired numerous follow-up methods including but not limited to CAGrad [25], PCGrad [37], GradDrop [8], FAMO [24] and FairGrad [3] with a convergence guarantee in the deterministic setting with full-gradient computations. The theoretical understanding of the convergence and complexity of stochastic MOO is not well-developed until very recently. [27] proposed stochastic multi-gradient (SMG) as a stochastic version of MGDA, and established its convergence guarantee. [40] analyzed the non-convergence issues of MGDA, CAGrad and PCGrad in the stochastic setting, and further proposed a convergent approach named CR-MOGM. More recently, [13] and [6] proposed single-loop stochastic MOO methods named MoCo and MoDo, and proved their convergence to an $\epsilon$-accurate Pareto stationary point while guaranteeing an $\epsilon$-level average CA distance [^1] CA distance means the distance between the updating direction and the CA direction. Its formal definition can be found in Section 2.4 over all iterations. [35] proposed a double-loop algorithm named SDMGrad that enables to obtain an unbiased stochastic multi-gradient via a double-sampling strategy. They established the convergence of SDMGrad with a guaranteed $\epsilon$-level CA distance in every iteration, which we call as iteration-wise CA distance.
@@
However, all existing works are limited by the standard $L$-smooth and bounded-gradient assumptions. Nevertheless, a recent study [38] indicates that such assumptions may not necessarily be true for the training of neural networks and an alternative $(L_{0},L_{1})$-smoothness condition was observed and studied, which assumes the Lipschitz constant to be linear in the gradient norm and the gradient norm to be potentially infinite. Furthermore, this phenomenon has been consistently observed in our experiments (e.g., see Figure 1). Interestingly, it has been widely observed that MGDA algorithms always converge even under such generalized smoothness conditions [32, 25, 35]. This naturally raises a thought-provoking question:
@@
This question remains open due to the following challenges. The analysis of existing MOO methods cannot be generalized to this $(L_{0},L_{1})$-smoothness directly due to the possible unbounded smoothness or gradient norm. In addition, all existing works [31, 38, 22, 21, 19, 11, 9, 39] in generalized smoothness are limited to the single task problems, which are fundamentally different from the MOO problems since even though each single task is generalized smooth, the linear combination of these tasks is not necessarily generalized smooth. In this paper, we provide an affirmative answer to this question. Our main contributions are summarized below.
@@ 1.1 Our Contributions
Weakest assumptions in MOO. In this paper, we investigate the generalized $\ell$-smooth assumption, where $\ell$ is a general non-decreasing function of gradient norm, and includes both the standard $L$-smooth and $(L_{0},L_{1})$-smooth assumptions as special cases. This assumption finds many applications, such as LSTM models [38], transformers [11], distributionally robust optimization [19] and higher-order polynomial functions [9]. In addition, we do not make any bounded-gradient assumption, which is required in previous analysis to ensure the bounded multi-gradient approximation. To the best of our knowledge, this is the first work to investigate generalized smoothness in MOO problems. Table 1 presents a detailed comparison of assumptions with existing analyses.
Convergence analysis. We first show that the vanilla MGDA method can converge to an $\epsilon$-accurate Pareto stationary point, while guaranteeing a small $\epsilon$-level average CA distance. A warm start process is introduced before the main loop of MGDA to achieve a more aggressive $\epsilon$-level iteration-wise CA distance. In the stochastic setting, we provide the convergence guarantee for MGDA with a double sampling, which was introduced by [35, 6] to obtain unbiased multi-gradient approximations. Furthermore, we analyze a computation- and memory-efficient variant of MGDA, named MGDA with fast approximation (MGDA-FA), which updates the objective weights $w$ using only forward passes of $F(\cdot)$ rather than gradient $\nabla F$, effectively reducing $O(K)$ time and space to $O(1)$ without hurting the performance guarantee.
Sample complexity comparison. To achieve an $\epsilon$-accurate Pareto stationary point and an $\epsilon$-level average CA distance, we show that MGDA require $\mathcal{O}(\epsilon^{-2})$ and $\mathcal{O}(\epsilon^{-4})$ samples in the deterministic and stochastic settings, respectively. Both of these complexities match with the existing best results. Furthermore, to achieve an $\epsilon$-level iteration-wise CA distance, MGDA with warm start requires an increased number of samples, at the order of $\mathcal{O}(\epsilon^{-11})$ and $\mathcal{O}(\epsilon^{-17})$ respectively, in both deterministic and stochastic scenarios, due to smaller step sizes and mini-batch data sampling. Typically, achieving an $\epsilon$-level iteration-wise CA distance results in much higher sample complexity, such as $\mathcal{O}(\epsilon^{-24})$ in [13], $\mathcal{O}(\epsilon^{-16})$ in [6] and $\mathcal{O}(\epsilon^{-12})$ in [35] for non-convex stochastic setting. Moreover, we show that MGDA-FA achieves the same performance guarantee as vanilla MGDA.
@@ 1.2 Related Works
Gradient-based multi-objective optimization. A variety of gradient manipulation techniques have emerged for simultaneous learning of multiple tasks. One prevalent category of methods adjusts the weights of various objectives according to factors such as uncertainty [20], gradient norm [7], and training complexity [17]. Methods based on MOO have garnered increased attention due to their systematic designs, enhanced training stability, and model-agnostic nature. For instance, [32] framed Multi-Task Learning (MTL) as a MOO problem and introduced an optimization method akin to MGDA [12]. Afterward, many MGDA-based methods have been proposed to mitigate gradient conflict with promising empirical performance. Among them, PCGrad [37] avoids conflict by projecting the gradient of each task on the norm plane of other tasks. GradDrop [8] randomly drops out conflicted gradients. CAGrad [25] adds a constraint on the update direction to be close to the average gradient. NashMTL [29] and FairGrad [3] formulated MTL as a bargaining game and a resource allocation problem, respectively. Theoretically, [13] proposed a provably convergent stochastic MOO method named MoCo based on an auxiliary tracking variable for gradient approximation. [6] characterized the trade-off among optimization, generalization, and conflict avoidance in MOO. [35] proposed and analyzed a stochastic MOO method named SDMGrad with a preference-oriented regularizer. However, all these works rely on the $L$-smoothness and bounded-gradient assumptions. In contrast, this paper focuses on the MOO problems with generalized $\ell$-smooth objectives.
Method Smoothness 11footnotemark: 1 Assumption22footnotemark: 2 Setting
CAGrad [25] (LS) (OD) Deterministic
PCGrad [37] (LS) (C) or (BC) Deterministic
SMG [27] (LS) (BG) Stochastic
CR-MOGM [40] (LS) (BF) and (BG) Stochastic
MoCo [13] (LS) (BF) and (BG) Stochastic
MoDo [6] (LS) (BG) Stochastic
SDMGrad [35] (LS) (BG) Stochastic
MGDA (Them. 1 and 3 ) (GS) N/A Deterministic
Stochastic MGDA (Them. 2 and 6) (GS) N/A Stochastic
@@
Generalized smoothness. The generalized $(L_{0},L_{1})$-smoothness was firstly proposed by [38], which was observed from extensive empirical experiments in training neural networks. A clipping algorithm was developed by [38] and the convergence rate was provided. Later, [19] analyzed the convergence of a normalized momentum method. The SPIDER algorithm was also applied to solve generalized smooth problems in [31, 9], where [9] studied a new notion of $\alpha$-symmetric generalized smoothness, which includes $(L_{0},L_{1})$-smoothness as a special case. Very recently, a new generalized $\ell$-smoothness condition was studied in [21, 22], which is the weakest smoothness condition and includes all the smoothness conditions discussed above. However, all the existing works on generalized smoothness are limited to single-task optimizations and the understanding of MOO is insufficient. This paper provides the first study of MOO under the generalized $\ell$-smoothness condition.
@@ 2.1 Generalized smoothness
The standard $L$-smoothness condition is widely investigated in existing optimization studies [15, 16], which assumes a function $f:\mathcal{X}\to\mathbb{R}$ to be $L$-smooth if there exists a bounded constant $L$ such that for any $x,y\in\mathcal{X}$, $\|\nabla f(x)-\nabla f(y)\|\leq L\|x-y\|.$ Nevertheless, recent studies show that in the training of neural networks such as LSTM models [38], transformers [11], distributionally robust optimization [19] and high-order polynomials functions [9], the standard $L$-smoothness assumption does not hold. Instead, a generalized $(L_{0},L_{1})$-smoothness assumption was observed and studied in the training of LSTM models in [38], which assumes that for any $x\in\mathcal{X}$, $\|\nabla^{2}f(x)\|\leq L_{0}+L_{1}\|\nabla f(x)\|.$ This assumption implies the Lipschitz constant is potentially unbounded and reduces to the $L$-smoothness if $L_{1}=0$. Later, a more generalized assumption was proposed and studied in [21]:
@@
(Generalized $\ell$-smoothness, Definition 1 in [21]). A real-valued differentiable function $f:\mathcal{X}\rightarrow\mathbb{R}$ is generalized $\ell$-smooth if $\|\nabla^{2}f(x)\|\leq\ell(\|\nabla f(x)\|)$ almost everywhere in $\mathcal{X}$, where :[0,+)($\ell:[0,+\infty)\rightarrow(0,+\infty)$ is a continuous non-decreasing function.
The $(L_{0},L_{1})$-smoothness is a special case of generalized $\ell$-smoothness, where $\ell(a)=L_{0}+L_{1}a$. There is another definition of generalized smooth, which is widely used and is equivalent to the $\ell$-smoothness:
@@
(($r,\ell$)-smoothness, Definition 2 in [21]). A real-valued differentiable function $f:\mathcal{X}\rightarrow\mathbb{R}$ is $(r,\ell)$-smooth if 1) for any $x\in\mathcal{X},B(x,r(\|\nabla f(x)\|))\in\mathcal{X}$, and 2) for any $x_{1},x_{2}\in B(x,r(\|\nabla f(x)\|)),\|\nabla f(x_{1})-\nabla f(x_{2})\|\leq\ell(\|\nabla f(x)\|)\|x_{1}-x_{2}\|$, where for continuous functions $r,\ell:[0,+\infty)\rightarrow(0,+\infty)$, $r$ is non-increasing, $\ell$ is non-decreasing and $B(x,R)$ is the Euclidean ball centered at $x$ with radius $R$.
In $B(x,r(\|\nabla f(x)\|))$, the function $f$ is also $L$-smooth where $L=\ell(\|\nabla f(x)\|)$. Proposition 3.2 in [21] shows that Definition 1 and Definition 2 are equivalent: An $(r,\ell)$-smooth function is $\ell$-smooth; and a generalized $\ell$-smooth function satisfying Assumption 1 is $(r,m)$-smooth where $m(u):=\ell(u+a)$ and $r(u):=a/m(u)$ for any $a>0$.

### v5

@@ 1 Introduction
There have been a variety of emerging applications of multi-objective optimization (MOO), such as online advertising (Ma et al., 2018), autonomous driving (Huang et al., 2019), and reinforcement learning (Thomas et al., 2021). Mathematically, the MOO problem takes the following formulation.
@@
where $K$ is the total number of objectives and $f_{k}(x)$ is the $k$-objective function given model parameters $x$. Under the stochastic setting, $f_{k}(x)=\mathbb{E}_{s}[f_{k}(x;s)]$, where $s$ denotes data sample. In MOO setting, we are interested in optimizing all the objectives simultaneously. However, this problem is challenging due to the gradient conflict that some objectives with larger gradients dominate the update direction at the sacrifice of significant performance degeneration on the less-fortune objectives with smaller gradients. Thus, a widely-adopted target is to find the Pareto stationary point $x$ that the performance of all objectives cannot be further improved without compromising some objectives. A variety of MOO-based methods have been proposed to mitigate this conflict and find a more balanced solution among all objectives. In particular, the multiple gradient descent algorithm (MGDA) (Désidéri, 2012) aims to find a conflict-avoidant (CA) update direction that maximizes the minimal improvement among all objectives and converges to a Pareto stationary point at which there is no common descent direction for all objective functions. This idea then inspired numerous follow-up methods including but not limited to CAGrad (Liu et al., 2021), PCGrad (Yu et al., 2020), GradDrop (Chen et al., 2020), FAMO (Liu et al., 2024) and FairGrad (Ban & Ji, 2024) with a convergence guarantee in the deterministic setting with full-gradient computations. The theoretical understanding of the convergence and complexity of stochastic MOO is not well-developed until very recently. Liu & Vicente (2021) proposed stochastic multi-gradient (SMG) as a stochastic version of MGDA, and established its convergence guarantee. Zhou et al. (2022) analyzed the non-convergence issues of MGDA, CAGrad and PCGrad in the stochastic setting, and further proposed a convergent approach named CR-MOGM. More recently, Fernando et al. (2022) and Chen et al. (2024) proposed single-loop stochastic MOO methods named MoCo and MoDo, and proved their convergence to an $\epsilon$-accurate Pareto stationary point while guaranteeing an $\epsilon$-level average CA distance [^1] CA distance means the distance between the updating direction and the CA direction. Its formal definition can be found in Section 2.4 over all iterations. Xiao et al. (2024) proposed a double-loop algorithm named SDMGrad that enables to obtain an unbiased stochastic multi-gradient via a double-sampling strategy. They established the convergence of SDMGrad with a guaranteed $\epsilon$-level CA distance in every iteration, which we call as iteration-wise CA distance.
@@
However, all existing works are limited by the standard $L$-smooth and bounded-gradient assumptions. Nevertheless, a recent study (Zhang et al., 2019) indicates that such assumptions may not necessarily be true for the training of neural networks and an alternative $(L_{0},L_{1})$-smoothness condition was observed and studied, which assumes the Lipschitz constant to be linear in the gradient norm and the gradient norm to be potentially infinite. Furthermore, this phenomenon has been consistently observed in our experiments (e.g., see Figure 1). Interestingly, it has been widely observed that MGDA algorithms always converge even under such generalized smoothness conditions (Sener & Koltun, 2018; Liu et al., 2021; Xiao et al., 2024). This naturally raises a thought-provoking question:
@@
This question remains open due to the following challenges. The analysis of existing MOO methods cannot be generalized to this $(L_{0},L_{1})$-smoothness directly due to the possible unbounded smoothness or gradient norm. In addition, all existing works (Reisizadeh et al., 2023; Zhang et al., 2019; Li et al., 2024b; a; Jin et al., 2021; Crawshaw et al., 2022; Chen et al., 2023; Zhang et al., 2024) in generalized smoothness are limited to the single task problems, which are fundamentally different from the MOO problems since even though each single task is generalized smooth, the linear combination of these tasks is not necessarily generalized smooth. In this paper, we provide an affirmative answer to this question. Our main contributions are summarized below.
@@ 1.1 Our Contributions
Weakest assumptions in MOO. In this paper, we investigate the generalized $\ell$-smooth assumption, where $\ell$ is a general non-decreasing function of gradient norm, and includes both the standard $L$-smooth and $(L_{0},L_{1})$-smooth assumptions as special cases. This assumption finds many applications, such as LSTM models (Zhang et al., 2019), transformers (Crawshaw et al., 2022), distributionally robust optimization (Jin et al., 2021) and higher-order polynomial functions (Chen et al., 2023). In addition, we do not make any bounded-gradient assumption, which is required in previous analysis to ensure the bounded multi-gradient approximation. To the best of our knowledge, this is the first work to investigate generalized smoothness in MOO problems. Table 1 presents a detailed comparison of assumptions with existing analyses.
Convergence analysis. We first show that the vanilla MGDA method can converge to an $\epsilon$-accurate Pareto stationary point, while guaranteeing a small $\epsilon$-level average CA distance. A warm start process is introduced before the main loop of MGDA to achieve a more aggressive $\epsilon$-level iteration-wise CA distance. In the stochastic setting, we provide the convergence guarantee for MGDA with a double sampling, which was introduced by Xiao et al. (2024); Chen et al. (2024) to obtain unbiased multi-gradient approximations. Furthermore, we analyze a computation- and memory-efficient variant of MGDA, named MGDA with fast approximation (MGDA-FA), which updates the objective weights $w$ using only forward passes of $F(\cdot)$ rather than gradient $\nabla F$, effectively reducing $O(K)$ time and space to $O(1)$ without hurting the performance guarantee.
Sample complexity comparison. To achieve an $\epsilon$-accurate Pareto stationary point and an $\epsilon$-level average CA distance, we show that MGDA require $\mathcal{O}(\epsilon^{-2})$ and $\mathcal{O}(\epsilon^{-4})$ samples in the deterministic and stochastic settings, respectively. Both of these complexities match with the existing best results. Furthermore, to achieve an $\epsilon$-level iteration-wise CA distance, MGDA with warm start requires an increased number of samples, at the order of $\mathcal{O}(\epsilon^{-11})$ and $\mathcal{O}(\epsilon^{-17})$ respectively, in both deterministic and stochastic scenarios, due to smaller step sizes and mini-batch data sampling, shown in Table 1. Typically, achieving an $\epsilon$-level iteration-wise CA distance results in much higher sample complexity, such as $\mathcal{O}(\epsilon^{-24})$ in Fernando et al. (2022), $\mathcal{O}(\epsilon^{-12})$  [^1] The order differs from (Chen et al., 2024) due to different definitions of the $\epsilon$-accurate Pareto stationarity and is taken when both $\epsilon$-accurate Pareto stationarity and $\epsilon$-level iteration-wise CA distance can be achieved. in Chen et al. (2024) and $\mathcal{O}(\epsilon^{-12})$ in Xiao et al. (2024) for non-convex stochastic setting. Moreover, we show that MGDA-FA achieves the same performance guarantee as vanilla MGDA.
@@ 1.2 Related Works
Gradient-based multi-objective optimization. A variety of gradient manipulation techniques have emerged for simultaneous learning of multiple tasks. One prevalent category of methods adjusts the weights of various objectives according to factors such as uncertainty (Kendall et al., 2018), gradient norm (Chen et al., 2018), and training complexity (Guo et al., 2018). Methods based on MOO have garnered increased attention due to their systematic designs, enhanced training stability, and model-agnostic nature. For instance, Sener & Koltun (2018) framed Multi-Task Learning (MTL) as a MOO problem and introduced an optimization method akin to MGDA (Désidéri, 2012). Afterward, many MGDA-based methods have been proposed to mitigate gradient conflict with promising empirical performance. Among them, PCGrad (Yu et al., 2020) avoids conflict by projecting the gradient of each task on the norm plane of other tasks. GradDrop (Chen et al., 2020) randomly drops out conflicted gradients. CAGrad (Liu et al., 2021) adds a constraint on the update direction to be close to the average gradient. NashMTL (Navon et al., 2022) and FairGrad (Ban & Ji, 2024) formulated MTL as a bargaining game and a resource allocation problem, respectively. Theoretically, Fernando et al. (2022) proposed a provably convergent stochastic MOO method named MoCo based on an auxiliary tracking variable for gradient approximation. Chen et al. (2024) characterized the trade-off among optimization, generalization, and conflict avoidance in MOO. Xiao et al. (2024) proposed and analyzed a stochastic MOO method named SDMGrad with a preference-oriented regularizer. However, all these works rely on the $L$-smoothness and bounded-gradient assumptions. In contrast, this paper focuses on the MOO problems with generalized $\ell$-smooth objectives.
Method Smoothness 11footnotemark: 1 Assumption22footnotemark: 2 Setting Complexity 33footnotemark: 3
CAGrad (Liu et al., 2021) (LS) (OD) Deterministic N/A
PCGrad (Yu et al., 2020) (LS) (C) or (BC) Deterministic N/A
SMG (Liu & Vicente, 2021) (LS) (C) or (BC) Stochastic N/A
CR-MOGM (Zhou et al., 2022) (LS) (BF) and (BG) Stochastic $\mathcal{O}(\epsilon^{-4})$
MoCo (Fernando et al., 2022) (LS) (BF) and (BG) Stochastic $\mathcal{O}(\epsilon^{-4})$
MoDo (Chen et al., 2024) (LS) (BG) Stochastic $\mathcal{O}(\epsilon^{-4})$
SDMGrad (Xiao et al., 2024) (LS) (BG) Stochastic $\mathcal{O}(\epsilon^{-4})$
MGDA (Them. 1 and 3 ) (GS) N/A Deterministic $\mathcal{O}(\epsilon^{-2})$
Stochastic MGDA (Them. 2 and 6) (GS) N/A Stochastic $\mathcal{O}(\epsilon^{-4})$
@@
Generalized smoothness. The generalized $(L_{0},L_{1})$-smoothness was firstly proposed by Zhang et al. (2019), which was observed from extensive empirical experiments in training neural networks. A clipping algorithm was developed by Zhang et al. (2019) and the convergence rate was provided. Later, Jin et al. (2021) analyzed the convergence of a normalized momentum method. The SPIDER algorithm was also applied to solve generalized smooth problems in Reisizadeh et al. (2023); Chen et al. (2023), where Chen et al. (2023) studied a new notion of $\alpha$-symmetric generalized smoothness, which includes $(L_{0},L_{1})$-smoothness as a special case. Very recently, a new generalized $\ell$-smoothness condition was studied in Li et al. (2024a; b), which is the weakest smoothness condition and includes all the smoothness conditions discussed above. However, all the existing works on generalized smoothness are limited to single-task optimizations and the understanding of MOO is insufficient. This paper provides the first study of MOO under the generalized $\ell$-smoothness condition.
@@ 2.1 Generalized smoothness
The standard $L$-smoothness condition is widely investigated in existing optimization studies (Ghadimi & Lan, 2013; Ghadimi et al., 2016), which assumes a function $f:\mathcal{X}\to\mathbb{R}$ to be $L$-smooth if there exists a bounded constant $L$ such that for any $x,y\in\mathcal{X}$, $\|\nabla f(x)-\nabla f(y)\|\leq L\|x-y\|.$ Nevertheless, recent studies show that in the training of neural networks such as LSTM models (Zhang et al., 2019), transformers (Crawshaw et al., 2022), distributionally robust optimization (Jin et al., 2021) and high-order polynomials functions (Chen et al., 2023), the standard $L$-smoothness assumption does not hold. Instead, a generalized $(L_{0},L_{1})$-smoothness assumption was observed and studied in the training of LSTM models in Zhang et al. (2019), which assumes that for any $x\in\mathcal{X}$, $\|\nabla^{2}f(x)\|\leq L_{0}+L_{1}\|\nabla f(x)\|.$ This assumption implies the Lipschitz constant is potentially unbounded and reduces to the $L$-smoothness if $L_{1}=0$. Later, a more generalized assumption was proposed and studied in Li et al. (2024a):
@@
(Generalized $\ell$-smoothness, Definition 1 in Li et al. (2024a)). A real-valued differentiable function $f:\mathcal{X}\rightarrow\mathbb{R}$ is generalized $\ell$-smooth if $\|\nabla^{2}f(x)\|\leq\ell(\|\nabla f(x)\|)$ almost everywhere in $\mathcal{X}$, where :[0,+)($\ell:[0,+\infty)\rightarrow(0,+\infty)$ is a continuous non-decreasing function.
The $(L_{0},L_{1})$-smoothness is a special case of generalized $\ell$-smoothness, where $\ell(a)=L_{0}+L_{1}a$. Another definition of generalized smooth is widely used and equivalent to the $\ell$-smoothness:
@@
(($r,\ell$)-smoothness, Definition 2 in Li et al. (2024a)). A real-valued differentiable function $f:\mathcal{X}\rightarrow\mathbb{R}$ is $(r,\ell)$-smooth if 1) for any $x\in\mathcal{X},B(x,r(\|\nabla f(x)\|))\in\mathcal{X}$, and 2) for any $x_{1},x_{2}\in B(x,r(\|\nabla f(x)\|)),\|\nabla f(x_{1})-\nabla f(x_{2})\|\leq\ell(\|\nabla f(x)\|)\|x_{1}-x_{2}\|$, where for continuous functions $r,\ell:[0,+\infty)\rightarrow(0,+\infty)$, $r$ is non-increasing, $\ell$ is non-decreasing and $B(x,R)$ is the Euclidean ball centered at $x$ with radius $R$.
In $B(x,r(\|\nabla f(x)\|))$, $f$ is also $L$-smooth where $L=\ell(\|\nabla f(x)\|)$. Definitions 1 and 2 are equivalent (Li et al., 2024a): An $(r,\ell)$-smooth function is $\ell$-smooth; and a $\ell$-smooth function satisfying Assumption 1 is $(r,m)$-smooth with $m(u):=\ell(u+a)$ and $r(u):=a/m(u)$ for any $a>0$.